OpenStudy (anonymous):

For the parent function y=√x what effect does a value of h = 5 have on the graph?

OpenStudy (anonymous):

typically, y=a*f(x-h)+k where a is a vertical stretch or shrink, h is a horizontal shift and k is a vertical shift. is that how it is presented in your class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if \(y=\sqrt{x-h}=\sqrt{x-5}\) do? Think about \(y=\sqrt{x}\) and what you can put ina and what you get out.

OpenStudy (anonymous):

ok can you keep solving it to completion soi can see

OpenStudy (anonymous):

what is the smallest x can be for the transformed function?